# Self-Admitted Technical Debt in LLM Software: An Empirical Comparison with ML and Non-ML Software

### v2

@@ VI Implications
For Practitioners: prevent debt early, especially in critical stages. The first SATD introduced in a file has a <<5% removal rate, making early architectural decisions effectively permanent, hence, developers should apply stricter reviews when introducing new files. Large commits (>>500 lines or >>10 files) predict 34x3–4x longer debt persistence, warranting CI/CD warnings. Debt concentrates in deployment, monitoring, and pretraining (RQ3), requiring focused refactoring. With 18% of LLM debt from external issues like model-stack workarounds (RQ2), teams need resources to track upstream changes and maintain compatibility layers.
For Tool Builders: paradigm-specific detection and prevention. RQ2 identified three new debt categories unique to LLM development; existing SATD detectors trained on traditional or pre-LLM ML comments will miss these patterns, requiring retraining on the extended taxonomy. RQ5 shows SATD prediction accuracy varies by paradigm (LLM: 0.82, ML: 0.96, non-ML: 0.74), necessitating paradigm-specific tooling. RQ3 suggests stage-aware warnings should prioritize debt in deployment and pretraining code over evaluation or alignment scripts.
@@
For the broader SE community: LLM development as a hybrid paradigm. Across all five research questions, LLM development emerges as a distinct hybrid paradigm. RQ1 shows comparable debt rates to ML (3.95% vs. 4.1%), yet RQ4 reveals different temporal patterns (2.4x longer debt-free periods followed by rapid accumulation). RQ2 shows debt arising from external ecosystem volatility; RQ3 reveals debt concentrates in operational infrastructure rather than model logic. This challenges assumptions in traditional software (stable dependencies, full control) and ML systems (experimental debt, rapid iteration). Converging debt levels despite differing temporal patterns suggests technical debt is fundamental to software evolution, not AI-specific.
@@ VII Threats to Validity
Construct Validity. Our analysis uses SATD as a proxy for technical debt, which excludes non-admitted, implicit, or systemic debt. This limits findings to explicitly documented debt, potentially underestimating total technical debt. The NLP-based detector achieved 0.82 F1 in prior evaluations but produced 44% false positives in our dataset, requiring manual filtering that may introduce selection bias. The taxonomy extension relies on 377 manually classified instances with 85.7% initial agreement; the 14.3% requiring discussion may reflect ambiguous debt boundaries unique to LLM systems. Broadening Design Debt for non-OO architectures may complicate comparisons with prior object-oriented work.
Internal Validity. Repository age confounds temporal comparisons: LLM repositories (post-2022) are younger than ML repositories (2015–2021), making it unclear whether differences reflect paradigm shifts or maturation effects. The 44x44xincrease in ML file-level introduction time (10 days in 2021 vs. 441 days in 2025) illustrates how repository age alters debt patterns, yet young LLM projects are compared against mature ML projects. Survival analysis censors unresolved SATD at study endpoints, potentially inflating persistence metrics for recent commits. Random Forest models may miss LLM-specific factors like prompt quality, API stability, or RAG complexity. Master-branch-only analysis excludes debt in experimental branches.
External Validity. Findings are limited to open-source Python projects on GitHub; generalizability to other languages, closed-source systems, or different contexts is uncertain. LLM repository selection focused on post-ChatGPT projects using specific frameworks (LangChain, LlamaIndex) and may not represent all LLM development patterns, particularly enterprise systems or research codebases for novel architectures. The 159-repository sample per category, while statistically powered for prevalence comparisons, may miss rare debt patterns in specialized domains (e.g., medical or financial LLMs). Pipeline stage mapping relies on manual classification of file paths and content, potentially misclassifying cross-cutting concerns or hybrid components. GitHub projects may overrepresent Western, English-speaking, startup contexts compared to industrial AI development elsewhere.
@@ VIII Conclusion
This study presents the first systematic replication and extension of SATD analysis in the LLM era. Applying Bhatia et al.’s methodology to 477 LLM, ML, and non-ML repositories reveals both continuity and evolution in technical debt patterns. Our contributions include: (1) evidence that SATD prevalence gaps across software categories are narrowing, (2) identification of three LLM-specific debt types extending existing taxonomies, (3) empirical demonstration of repository maturation effects, and (4) a reproducible framework for longitudinal SATD analysis. SATD pattern convergence across paradigms suggests technical debt is fundamental to software evolution, not a sign of poor practice. As LLM systems become critical infrastructure, understanding their debt accumulation is essential for sustainable AI development. Future work should explore CI/CD bots to flag high-risk pull requests and prevent lasting debt. Researchers should expand the SATD taxonomy with larger datasets or domain-specific LLM applications to uncover additional debt patterns. Qualitative studies with LLM practitioners could reveal how debt tradeoffs are made and challenges of resolving foundational debt.

### v3

@@ VI Implications
For Practitioners: prevent debt early, especially in critical stages. The first SATD introduced in a file has a <<5% removal rate, making early architectural decisions effectively permanent; developers should therefore enforce stricter reviews for new files. Large commits (>>500 lines or >>10 files) correlate with 343–4 times longer debt persistence and should trigger CI/CD warnings. Debt concentrates in deployment, monitoring, and pretraining (RQ3), warranting targeted refactoring. Since 22.4% of LLM debt comes from external issues like model-stack workarounds (RQ2), teams should track upstream changes and maintain compatibility layers.
For Tool Builders: paradigm-specific detection and prevention. RQ2 identifies three LLM-specific debt categories that existing SATD detectors, trained on traditional ML or non-ML comments, will miss. Therefore, detectors should be retrained on LLM-specific taxonomies. RQ5 shows SATD prediction accuracy differs by paradigm (LLM: 0.82, ML: 0.96, non-ML: 0.74), motivating paradigm-specific tooling. RQ3 suggests stage-aware warnings should prioritize debt in deployment and pretraining code over evaluation or alignment scripts.
@@
For the broader SE community: LLM development as a hybrid paradigm. Across all five research questions, LLM development emerges as a distinct hybrid paradigm. RQ1 shows comparable debt rates to ML (3.95% vs. 4.10%), but RQ4 reveals different timing (2.4x longer debt-free periods followed by rapid accumulation). RQ2 attributes debt to external ecosystem volatility; RQ3 shows it concentrates in operational infrastructure rather than model logic. Thus, LLM development differs from both traditional software engineering (which assumes stable dependencies) and classical ML (which assumes experimental workflows). Converging debt levels despite differing temporal patterns suggests technical debt is fundamental to software evolution, not AI-specific.
@@ VII Threats to Validity
Construct Validity. Our analysis treats SATD as a proxy for technical debt, excluding non-admitted, implicit, and systemic debt, and thus likely underestimates total debt. The NLP-based detector, previously reporting 0.82 F1-score, yielded 44% false positives in our dataset, requiring manual filtering that may introduce selection bias. The taxonomy extension relies on 377 manually classified instances with 85.7% initial agreement; the remaining 14.3% suggest ambiguous, LLM-specific debt boundaries. Expanding Design Debt beyond OO architectures may hinder comparison with prior OO-focused studies.
Internal Validity. Repository age confounds temporal comparisons: LLM repositories (post-2022) are younger than ML repositories (2015–2021), making it unclear whether differences reflect age or paradigm shifts. The $44x$ increase in ML file-level introduction time (10 days in 2021 vs. 441 days in 2025) shows how age alters debt patterns, yet young LLM projects are compared with mature ML projects. Survival analysis censors unresolved SATD at study endpoints, likely inflating persistence metrics for recent commits. Random Forest models may overlook LLM-specific factors such as prompt quality, API stability, or RAG complexity. Focusing only on the master branch omits debt in experimental branches.
External Validity. Findings are limited to open-source Python projects on GitHub; generalizability to other languages, closed-source systems, or different contexts is uncertain. LLM repository selection focused on post-ChatGPT projects using specific frameworks (LangChain, LlamaIndex) and may not represent all LLM development patterns, particularly enterprise systems or research codebases developing novel architectures. The 159-repository sample per category, while statistically powered for prevalence comparisons, may miss rare debt patterns in specialized domains (e.g., medical or financial LLMs). Pipeline stage mapping relies on manual classification of file paths and content, potentially misclassifying cross-cutting concerns or hybrid components. GitHub projects may overrepresent Western, English-speaking, startup contexts compared to industrial AI development elsewhere.
@@ VIII Conclusion
This study presents the first systematic replication and extension of SATD analysis in the LLM era. Applying Bhatia et al.’s methodology to 477 LLM, ML, and non-ML repositories reveals both continuity and evolution in technical debt patterns. Our contributions include: (1) evidence that SATD prevalence gaps across software categories are narrowing, (2) identification of three LLM-specific debt types extending existing taxonomies, (3) analysis of repository maturation effects, and (4) a reproducible framework for longitudinal SATD analysis. SATD pattern convergence across paradigms suggests technical debt is inherent to software evolution, not a sign of poor practice. Understanding LLM-specific debt accumulation is essential for sustainable AI development. Future work should develop CI/CD bots to flag high-risk pull requests and prevent lasting debt. Researchers should extend the SATD taxonomy using larger datasets or domain-specific LLMs to uncover additional debt patterns. Qualitative studies with LLM practitioners could clarify how they weigh debt tradeoffs and the challenges of resolving foundational debt.
Acknowledgment
This work is funded by the Natural Sciences and Engineering Research Council of Canada (NSERC): RGPIN-2023-05440 and RGPIN-2025-05897.